Circular No. 8688

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2006 E1 R. H. McNaught reports his discovery of a comet with a coma of di-

ameter ∼ 6′′, extended to the northwest, on Siding Spring Survey images taken with the 0.5-m Uppsala Schmidt telescope.

2006 UT  α_{2000}  δ_{2000}  Mag.
Mar. 11.74071  19^{h}31^{m}31.^{s}33  −46^{o}07′56′′.2  18.0
11.74600  19 31 31.51  −46 07 58.2  18.0
11.75130  19 31 31.72  −46 08 00.4  18.3
11.75661  19 31 31.94  −46 08 02.9  18.1
11.78456  19 31 32.97  −46 08 14.5
11.78789  19 31 33.09  −46 08 16.1  18.5
11.79123  19 31 33.16  −46 08 17.1  18.7
12.74392  19 32 08.04  −46 14 53.0  18.4
12.75994  19 32 08.68  −46 15 00.2  17.9

RS OPHIUCHI T. J. O’Brien, T. W. B. Muxlow, S. T. Garrington, and R. J. Davis,

Jodrell Bank Observatory, University of Manchester; R. W. Porcas, Max-Planck-Institut für Radioastronomie, Bonn; M. F. Bode, Liverpool John Moores University; S. P. S. Eyres, University of Central Lancashire; and A. Evans, Keele University, write that 6-cm radio observations made with the Very Long Baseline Array on Feb. 26 show radio emission from RS Oph in the form of a ring of diameter ≈ 18 mas, consistent with the size of the single circular component fitted to the MERLIN imaging reported on IAUC 8684. The ring appears almost circular, clumpy, significantly brighter on its eastern side, and with some evidence for a central point source. At a distance of 1600 pc, the diameter of the ring is 29 AU and, assuming that it has expanded uniformly since outburst, its expansion velocity is about 1800 km/s. O’Brien et al. suggest that this ring of radio emission may be identified with a shock wave expanding through the red-giant wind following the nova explosion, also resulting in the x-ray emission reported on IAUCs 8675, 8677, and 8683. VLBA observations are continuing.

Visual magnitude estimates, reported in part by E. Waagen, AAVSO: Feb. 17.249 UT, 6.5 (C. Labordena, Castellón, Spain); 21.847, 7.4 (A. Pearce, Nedlands, W. Australia); 26.180, 7.9 (B. Granslo, Fjellhamar, Norway); Mar. 2.163, 8.1 (C. Otten, Kinrooi, Belgium); 5.139, 8.2 (G. Mavrofridis, Nikea, Greece); 13.183, 8.9 (E. Muyllaert, Oostende, Belgium).

2006 March 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green